OpenStudy (anonymous):

190p+10f=850 & 180p+20f=900 f=calories in a gram of fat P=calories in a gram of protein

OpenStudy (anonymous):

i was trying 2 solve it my self but when i wanted to verify the answer its wrong i just need u to help me plz

OpenStudy (anonymous):

\[180p+20f=900 \]

OpenStudy (anonymous):

then i diveded 400/280

OpenStudy (anonymous):

o okies are supposed to completely cancel out a variable it does not work like regular single variable substitution

OpenStudy (anonymous):

x - y = 0 cancel out any variable 'x' or 'y' x - y = 0 im going to choose 'x' x - y = 0 <---------------------------------| + x - y = 0 cancel out use multiply by additive inverse of the 'x' value here ----------- x - y = 0 + x - y = 0 ------------ y = 0 y = value is x - y = 0 <---------substitute x using in this formula x - y = 0 <--- if you used this formula for elimination to isolate 'y' x - y = 0 <--- if you used this formula to isolate 'y' x - y = 0 <----------- substitute 'x' using this formula x= y=

OpenStudy (anonymous):

you can simplify it a bit Factor them out [ 190p+10f=850 ] [ 180p+20f=900 ] 19 p + f = 85 9 p + f = 45

OpenStudy (anonymous):

that should be much easier to look at and easier to solve